# Model Reconstruction Using Counterfactual Explanations: Mitigating the Decision Boundary Shift

### v1

@@ B.1 Details of Real-World Datasets
Comparing performance over four real-world dataset
@@
Fig. 11 illustrates the fidelities achieved by the two model architectures described above. Fig. 12 shows the corresponding variances of the fidelity values over 100 realizations. It can be observed that the variances diminish as the query size grows, indicating more stable model extractions. Fig. 13 demonstrates the effect of the proposed loss function in mitigating the decision boundary shift issue.
Figure 11: Fidelity for real-world datasets. Solid lines indicate “CCA” models. Dashed lines indicate “Baseline” models. Colors correspond to the model architecture, i.e., red: model 0, blue: model 1.
Figure 12: Variance of fidelity for real-world datasets. Solid lines indicate “CCA” models. Dashed lines indicate “Baseline” models. Colors correspond to the model architecture, i.e., red: model 0, blue: model 1.
@@
Empirical and theoretical rates of convergence
Fig. 14 compares the rate of convergence of the empirical approximation error i.e., $1-\mathbb{E}\left[{{\rm Fid}_{m,{\mathbb{D}}_{\text{ref}}}(\tilde{M}_{n})}\right]$ for two of the above experiments with the rate predicted by Theorem A.3. Notice how the empirical error decays faster than $n^{-2/(d-1)}$.
@@
Studying effects of Lipschitz constants
For this experiment, we use a target model having 3 hidden layers with the architecture (20, 10, 5) and a surrogate model having 2 hidden layers with the architecture (20, 10). The surrogate model layers are $L_{2}$-regularized with a fixed regularization coefficient of 0.001. We achieve different Lipschitz constants for the target models by controlling their $L_{2}$-regularization coefficients during the target model training step. Following (Gouk et al. 2021), we approximate the Lipschitz constant of target models by the product of the spectral norms of the weight matrices.
@@
Studying different model architectures
We use the HELOC dataset for this experiment. The target model has a fixed architecture of (20, 30, 10). Surrogate model architectures are given in Table 2, ranging from very complex to much simpler than the target model. The ensemble size $S$ is selected to be 50.
Table 2: Surrogate model architectures
Model Hidden layers
Model 0 (30, 20, 20, 10)
Model 1 (20, 30, 10)
Model 2 (20, 10)
Model 3 (5, 5)
Results (see Fig. 15) indicate that the performance is weakly dependent on the model architecture as long as the surrogate models are closer to the target model in complexity, which is an observation made also by Aïvodji, Bolot, and Gambs (2020). Results of additional experiments with varying target and surrogate model architectures are presented in Table 3 and 4.
Figure 15: Effect of the surrogate model architecture on the attack performance. Dashed lines correspond to “Baseline” models and solid lines represent the “CCA” models. See Table 2 for the model specifications. The attack performance is weakly dependent on the surrogate model architecture as long as it is closer to the complexity of the target model.
Table 3: Fidelity over ${\mathbb{D}}_{\text{test}}$ and ${\mathbb{D}}_{\text{uni}}$ for datasets DCCC and HELOC
Dataset: DCCC - Fidelity over ${\mathbb{D}}_{\text{test}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.96 0.98 0.97 0.99 0.96 0.99 0.97 0.99 0.91 0.97 0.92 0.98
(20,10,5) 0.96 0.98 0.97 0.99 0.94 0.97 0.96 0.98 0.91 0.97 0.92 0.97
(20,20,10,5) 0.93 0.97 0.94 0.98 0.92 0.97 0.93 0.97 0.88 0.94 0.92 0.96
Dataset: DCCC - Fidelity over ${\mathbb{D}}_{\text{uni}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.95 0.96 0.96 0.97 0.97 0.97 0.97 0.98 0.98 0.98 0.98 0.98
(20,10,5) 0.94 0.96 0.95 0.96 0.95 0.94 0.94 0.96 0.97 0.98 0.98 0.98
(20,20,10,5) 0.92 0.93 0.95 0.94 0.96 0.94 0.95 0.96 0.94 0.92 0.96 0.94
Dataset: HELOC - Fidelity over ${\mathbb{D}}_{\text{test}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
@@
Dataset: HELOC - Fidelity over ${\mathbb{D}}_{\text{uni}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
@@
Table 4: Fidelity over ${\mathbb{D}}_{\text{test}}$ and ${\mathbb{D}}_{\text{uni}}$ for datasets Adult Income and COMPAS
Dataset: Adult Income - Fidelity over ${\mathbb{D}}_{\text{test}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.88 0.89 0.92 0.93 0.82 0.84 0.91 0.93 0.94 0.95 0.95 0.96
(20,10,5) 0.87 0.88 0.91 0.93 0.79 0.82 0.90 0.92 0.93 0.94 0.95 0.96
Dataset: Adult Income - Fidelity over ${\mathbb{D}}_{\text{uni}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.71 0.81 0.75 0.87 0.78 0.84 0.79 0.87 0.84 0.88 0.85 0.91
(20,10,5) 0.71 0.78 0.74 0.83 0.77 0.82 0.78 0.85 0.82 0.88 0.84 0.90
Dataset: COMPAS - Fidelity over ${\mathbb{D}}_{\text{test}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.93 0.96 0.94 0.97 0.92 0.94 0.94 0.96 0.94 0.96 0.95 0.97
(20,10,5) 0.92 0.95 0.94 0.97 0.92 0.93 0.95 0.95 0.94 0.96 0.95 0.97
Dataset: COMPAS - Fidelity over ${\mathbb{D}}_{\text{uni}}$
Target archi. $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
100 queries 200 queries 100 queries 200 queries 100 queries 200 queries
Surrogate archi. $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.94 0.95 0.94 0.96 0.95 0.95 0.95 0.96 0.96 0.95 0.96 0.96
(20,10,5) 0.93 0.95 0.94 0.95 0.94 0.92 0.95 0.92 0.95 0.96 0.96 0.96
Studying alternate counterfactual generating method
@@
Counterfactuals can be generated such that they satisfy additional desirable properties such as actionability, sparsity and closeness to the data manifold, other than the proximity to the original instance. In this experiment, we observe how counterfactuals with above properties affect the attack performance. DCCC is used as the dataset. Target model has the architecture (20, 30, 10) and the architecture of the surrogate model is (10, 20).
@@
To generate actionable counterfactuals, we use Diverse Counterfactual Explanations (DiCE)by Mothilal, Sharma, and Tan (2020) with features “SEX”, “MARRIAGE”, and “AGE” kept unchanged. The diversity factor of DiCE generator is set to 1 in order to obtain only a single counterfactual for each query. Sparse counterfactuals are obtained by the same MCCC generator used in other experiments, but now with $L_{1}$ norm as the cost function $c(\boldsymbol{x},\boldsymbol{w})$. Counterfactuals from the data manifold (i.e., realistic counterfactuals, denoted by 1-NN) are generated using a 1-Nearest-Neighbour algorithm. Table 5 summarizes the performance of the attack.
When the counterfactuals lie further away from the decision boundary, the advantage of the proposed attack w.r.t. the baseline diminishes. In order to observe this effect, we carry-out the attacks using counterfactuals with two different levels of sparsity. It can be observed that when the sparsity increases, the counterfactuals tend to lie further away from the decision boundary (see Fig. 16). Accordingly, the performance gap between the CCA and the baseline attacks diminishes, as seen in Fig. 17.
Figure 16: Histogram of target model predictions $m(\boldsymbol{x})$ for instances with $\lfloor m(\boldsymbol{x})\rceil=0,1$ and counterfactuals, when different levels of sparsity is used with DiCE. Higher the parameter ’posthoc_sparsity_param’, more sparse the counterfactuals are. Note how the predictions on counterfactuals are concentrated away from the decision boundary when the sparsity is high. Dataset used is DCCC.
Figure 17: Attack performance on DCCC dataset when different levels of sparsity in DiCE is used. Higher the parameter ’posthoc_sparsity_param’, more sparse the counterfactuals are.
Table 5: Fidelity achieved with different counterfactual generating methods on DCCC dataset. Target model has hidden layers with neurons (20, 30, 10). Surrogate model architecture is (10, 20).
Fidelity over ${\mathbb{D}}_{\text{test}}$ Fidelity over ${\mathbb{D}}_{\text{uni}}$
CF method 100 queries 200 queries 100 queries 200 queries
Base. CCA Base. CCA Base. CCA Base. CCA
MCCF L2-norm 0.98 0.99 0.98 0.99 0.95 0.96 0.95 0.97
MCCF L1-norm 0.79 0.95 0.85 0.97 0.89 0.92 0.93 0.94
DiCE Actionable 0.98 0.98 0.98 0.98 0.92 0.94 0.91 0.94
1-NN 0.98 0.99 0.99 0.99 0.94 0.95 0.95 0.96
Studying effect of unbalanced ${\mathbb{D}}_{\text{attack}}$
In all the other experiments, the attack dataset ${\mathbb{D}}_{\text{attack}}$ used by the adversary is sampled from a class-wise balanced dataset. In this experiment we explore the effect of querying using an unbalanced ${\mathbb{D}}_{\text{attack}}$. Model architectures used are (20, 10) for the target model and surrogate model 0, and (20, 10, 5) for surrogate model 1.
Two versions of the Adult Income dataset were used in the experiment. The first version, used in Fig. 18(a) contains 24720 and 7841 examples from classes $y=0$ and $y=1$, respectively. The second version corresponding to Fig. 18(b) contains 2472 and 7841 examples from classes $y=0$ and $y=1$, respectively. The proposed attack surpasses the baseline attack in both scenarios.
Figure 18: Results corresponding to the Adult Income dataset with queries sampled from biased versions of the dataset (i.e., a biased ${\mathbb{D}}_{\text{attack}})$. The version used in Fig. 18(a) contains 24720 and 7841 examples from classes $y=0$ and $y=1$, respectively. The version corresponding to Fig. 18(b) contains 2472 and 7841 examples from classes $y=0$ and $y=1$, respectively.
Studying other machine learning models
@@
Studying alternate loss functions
@@
Figure 20: Performance of binary cross-entropy loss with labels 0, 0.5 and 1. Dashed lines corresponding to binary cross entropy (BCE) loss and solid lines depict the performance of the CCA loss.

### v2

@@ B.2.1 Visualizing the attack using synthetic data
B.2.2 Comparing performance over four real-world dataset
@@
Fig. 11 illustrates the fidelities achieved by the two model architectures described above. Fig. 12 shows the corresponding variances of the fidelity values over 100 realizations. It can be observed that the variances diminish as the query size grows, indicating more stable model reconstructions. Fig. 13 demonstrates the effect of the proposed loss function in mitigating the decision boundary shift issue.
Figure 11: Fidelity for real-world datasets. Blue lines indicate “CCA” models. Black lines indicate “Baseline” models.
Figure 12: Variance of fidelity for real-world datasets. Blue lines indicate “CCA” models. Black lines indicate “Baseline” models.
@@
B.2.3 Empirical and theoretical rates of convergence
Fig. 14 compares the rate of convergence of the empirical approximation error i.e., $1-\mathbb{E}\left[{{\rm Fid}_{m,{\mathbb{D}}_{\text{ref}}}(\tilde{M}_{n})}\right]$ for two of the above experiments with the rate predicted by Theorem 3.2. Notice how the empirical error decays faster than $n^{-2/(d-1)}$.
@@ B.2.3 Empirical and theoretical rates of convergence
B.2.4 Studying effects of Lipschitz constants
For this experiment, we use a target model having 3 hidden layers with the architecture (20, 10, 5) and a surrogate model having 2 hidden layers with the architecture (20, 10). The surrogate model layers are $L_{2}$-regularized with a fixed regularization coefficient of 0.001. We achieve different Lipschitz constants for the target models by controlling their $L_{2}$-regularization coefficients during the target model training step. Following Gouk et al. [2021], we approximate the Lipschitz constant of target models by the product of the spectral norms of the weight matrices.
@@ B.2.4 Studying effects of Lipschitz constants
B.2.5 Studying different model architectures
We observe the effect of the model architectures on the attack performance over Adult Income, COMPAS and HELOC datasets. Tables 3, 4, and 5, respectively, present the results.
Table 3: Fidelity over ${\mathbb{D}}_{\text{test}}$ and ${\mathbb{D}}_{\text{uni}}$ for Adult Income dataset
Target $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
${\mathbb{D}}_{\text{test}}$ n=100 n=200 n=100 n=200 n=100 n=200
Surrogate $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.88 0.89 0.92 0.93 0.82 0.84 0.91 0.93 0.94 0.95 0.95 0.96
(20,10,5) 0.87 0.88 0.91 0.93 0.79 0.82 0.90 0.92 0.93 0.94 0.95 0.96
Target $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
${\mathbb{D}}_{\text{uni}}$ n=100 n=200 n=100 n=200 n=100 n=200
Surrogate $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.71 0.81 0.75 0.87 0.78 0.84 0.79 0.87 0.84 0.88 0.85 0.91
(20,10,5) 0.71 0.78 0.74 0.83 0.77 0.82 0.78 0.85 0.82 0.88 0.84 0.90
Table 4: Fidelity over ${\mathbb{D}}_{\text{test}}$ and ${\mathbb{D}}_{\text{uni}}$ for COMPAS dataset
Target $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
${\mathbb{D}}_{\text{test}}$ n=100 n=200 n=100 n=200 n=100 n=200
Surrogate $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.93 0.96 0.94 0.97 0.92 0.94 0.94 0.96 0.94 0.96 0.95 0.97
(20,10,5) 0.92 0.95 0.94 0.97 0.92 0.93 0.95 0.95 0.94 0.96 0.95 0.97
Target $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
${\mathbb{D}}_{\text{uni}}$ n=100 n=200 n=100 n=200 n=100 n=200
Surrogate $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
(20,10) 0.94 0.95 0.94 0.96 0.95 0.95 0.95 0.96 0.96 0.95 0.96 0.96
(20,10,5) 0.93 0.95 0.94 0.95 0.94 0.92 0.95 0.92 0.95 0.96 0.96 0.96
Table 5: Fidelity over ${\mathbb{D}}_{\text{test}}$ and ${\mathbb{D}}_{\text{uni}}$ for HELOC dataset
Target $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
${\mathbb{D}}_{\text{test}}$ n=100 n=200 n=100 n=200 n=100 n=200
Surrogate $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
@@ B.2.5 Studying different model architectures
Target $\rightarrow$ (20,10) (20,10,5) (20,20,10,5)
${\mathbb{D}}_{\text{uni}}$ n=100 n=200 n=100 n=200 n=100 n=200
Surrogate $\downarrow$ Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA Base. CCA
@@
B.2.6 Studying alternate counterfactual generating method
@@
Counterfactuals can be generated such that they satisfy additional desirable properties such as actionability, sparsity and closeness to the data manifold, other than the proximity to the original instance. In this experiment, we observe how counterfactuals with above properties affect the attack performance. HELOC is used as the dataset. Target model has the architecture (20, 30, 10) and the architecture of the surrogate model is (10, 20).
@@
To generate actionable counterfactuals, we use Diverse Counterfactual Explanations (DiCE)by Mothilal et al. [2020] with the first four features, i.e., “estimate_of_risk”, “months_since_last_trade”, “average_duration_of_resolution”, and “number_of_satisfactory_trades” kept unchanged. The diversity factor of DiCE generator is set to 1 in order to obtain only a single counterfactual for each query. Sparse counterfactuals are obtained by the same MCCF generator used in other experiments, but now with $L_{1}$ norm as the cost function $c(\boldsymbol{x},\boldsymbol{w})$. Counterfactuals from the data manifold (i.e., realistic counterfactuals, denoted by 1-NN) are generated using a 1-Nearest-Neighbor algorithm. We use ROAR [Upadhyay et al., 2021] and C-CHVAE [Pawelczyk et al., 2020] to generate robust counterfactuals. Table 2 summarizes the performance of the attack. Fig. 8 shows the distribution of the counterfactuals generated using each method w.r.t. the decision boundary of the target model. We observe that the sparse, realistic, and robust counterfactuals have a tendency to lie farther away from the decision boundary, within the favorable region, when compared to the closest counterfactuals under $L_{2}$ norm.
B.2.7 Comparison with DualCFX Wang et al. [2022]
Wang et al. [2022] is one of the few pioneering works studying the effects of counterfactuals on model extraction, which proposes the interesting idea of using counterfactuals of counterfactuals to mitigate the decision boundary shift. This requires the API to provide counterfactuals for queries originating from both sides of the decision boundary. However, the primary focus of our work is on the one-sided scenario where an institution might be giving counterfactuals only to the rejected applicants to help them get accepted, but not to the accepted ones. Hence, a fair comparison cannot be achieved between CCA and the strategy proposed in Wang et al. [2022] in the scenario where only one-sided counterfactuals are available.
Therefore, in the two-sided scenario, we compare the performance of CCA with the DualCFX strategy proposed in Wang et al. [2022] under two settings:
  1. 1.
    only one sided counterfactuals are available for CCA (named CCA1)
  2. 2.
    CCA has all the data that DualCFX has (named CCA2)
We also include another baseline (following Aïvodji et al. [2020]) for the two-sided scenario where the models are trained only on query instances and counterfactuals, but not the counterfactuals of the counterfactuals. Results are presented in Table 6. Note that even for the same number of initial query instances, the total number of actual training instances change with the strategy being used (CCA1 < Baseline < DualCFX = CCA2 – e.g.: queries+CFs for the baseline but queries+CFs+CCFs for DualCFX).
Table 6: Comparison with DualCFX. Legend: Base.=Baseline model based on [Aïvodji et al., 2020], Dual=DualCFX, CCA1=CCA with one-sided counterfactuals, CCA2=CCA with counterfactuals from both sides.
Architecture known (model 0)
Dataset Query size ${\mathbb{D}}_{\text{test}}$ ${\mathbb{D}}_{\text{uni}}$
Base. Dual. CCA1 CCA2 Base. Dual. CCA1 CCA2
DCCC n=100 0.95 0.99 0.94 0.99 0.90 0.95 0.92 0.97
n=200 0.96 0.99 0.98 0.99 0.90 0.96 0.95 0.98
HELOC n=100 0.94 0.97 0.90 0.98 0.91 0.98 0.84 0.98
n=200 0.96 0.98 0.92 0.98 0.93 0.98 0.89 0.99
Architecture unknown (model 1)
Dataset Query size ${\mathbb{D}}_{\text{test}}$ ${\mathbb{D}}_{\text{uni}}$
Base. Dual. CCA1 CCA2 Base. Dual. CCA1 CCA2
DCCC n=100 0.92 0.98 0.93 0.98 0.88 0.92 0.89 0.93
n=200 0.96 0.99 0.96 0.99 0.89 0.94 0.94 0.96
HELOC n=100 0.92 0.91 0.90 0.96 0.88 0.92 0.84 0.96
n=200 0.95 0.92 0.91 0.97 0.93 0.94 0.88 0.97
B.2.8 Studying other machine learning models
@@
B.2.9 Studying effect of unbalanced ${\mathbb{D}}_{\text{attack}}$
In all the other experiments, the attack dataset ${\mathbb{D}}_{\text{attack}}$ used by the adversary is sampled from a class-wise balanced dataset. In this experiment we explore the effect of querying using an unbalanced ${\mathbb{D}}_{\text{attack}}$. Model architectures used are (20, 10) for the target model and surrogate model 0, and (20, 10, 5) for surrogate model 1. While the training set of the teacher and the test set of both the teacher and the surrogates were kept constant, the proportion of the samples in the attack set $\mathbb{D}_{\text{attack}}$ was changed. In the first case, examples from class $y=1$ were dominant (80%) and in the second case, the majority of the examples were from class $y=0$ (80%). The results are shown in Fig. 17.
Figure 17: Results corresponding to the HELOC dataset with queries sampled from biased versions of the dataset (i.e., a biased ${\mathbb{D}}_{\text{attack}})$. The version on the left uses a ${\mathbb{D}}_{\text{attack}}$ with 20% and 80% examples from classes $y=0$ and $y=1$, respectively. The version on the right was obtained with a ${\mathbb{D}}_{\text{attack}}$ comprising of 80% and 20% examples from classes $y=0$ and $y=1$, respectively.
B.2.10 Studying alternate loss functions
@@
Figure 18: Performance of binary cross-entropy loss with labels 0, 0.5 and 1. Black lines corresponding to binary cross entropy (BCE) loss and blue lines depict the performance of the CCA loss.